# The Binary and Ternary Quantization Can Improve Feature Discrimination

### v1

@@ 4.2.1 Setting
To validate the broad applicability of our feature discrimination analysis results, we here examine not only binary classification but also multiclass classification. Regarding the classifiers, we employ not only the fundamental KNN and SVM which mainly involve linear operations, but also more complicated classifiers involving nonlinear operations, such as the multilayer perceptron (MLP) (Rumelhart et al., 1986) and decision trees (Quinlan, 1986). For brevity, in the main text we present the classification results of YaleB, Newsgroup, and TIMIT using KNN (with Euclidean distance) and SVM, as illustrated in Figures 4 to 6. The results for other datasets, such as CIFAR10 and ImageNet1000, and other classifiers, including KNN with cosine distance, MLP and decision trees, are provided in Appendix C, specifically in Figures 14 to 20.
@@ 4.2.2 Results
Binary classification.
In Figures 4-6,14, 15 and 19, we present the binary classification results of KNN and SVM on five distinct datasets. It can be seen that there indeed exist specific ranges of quantization threshold $\tau$ values that allow both binary and ternary quantization to achieve superior, or comparable classification performance compared to the original data. Note that as illustrated in Figure 16, each data category approximately follows, rather than perfectly conforms to the Gaussian distribution assumption made in our theoretical analysis. This underscores the robustness of our theoretical results. Similar to the classification of synthetic data, the classification of real data exhibits the following performance trends. 1) When using Euclidean distance, KNN consistently obtains quantization threshold $\tau$ values that improve the classification accuracy of original data across all datasets. 2) In contrast to cosine distance, Euclidean distance enables KNN to cover a wider range of $\tau$ values that facilitate classification improvement. 3) With SVM, quantization occasionally achieves comparable performance to the original data, rather than surpassing it, as exemplified in Figure 14. 4) Ternary quantization often outperforms binary quantization by offering a broader range of threshold $\tau$ values conductive to classification improvement. The reasoning behind these results has been elaborated in our previous classification analysis of synthetic data. Overall, the consistency in classification performance between real data and synthetic data substantiates the rationality and generalizability of our theoretical findings.
@@
In addition to KNN and SVM, which employ linear operations to measure feature similarity and are thus well-suited for our linear discrimination analysis, more complex nonlinear classifiers, like MLP and decision trees, can also obtain the threshold $\tau$ values that improve classification, as demonstrated in Figures 17 and 18. This capability stems from the fact that these nonlinear classifiers are grounded in linear operations, which serve to measure the linear similarity between features or model parameters, thereby aligning with our analytical framework.
@@
By improving feature discrimination between two classes, the performance of multiclass classification is expected to improve as well. This improvement has been observed in recent experiments (Lu et al., 2023), particularly in the challenging 1,000-class classification on deep features of ImageNet1000, as illustrated in Figure 20. In multiclass classification, we simply apply a uniform threshold to each dimension of the feature vectors. This approach conforms to our analytical framework rooted in binary classification, as classification in each dimension can be approximately considered a binary classification problem, where the values of feature elements across different classes indicate the intensity of the feature attribute at that dimension, i.e., whether it is strong or weak. The successful adaption to multiclass classification highlights the broad applicability of our theoretical findings.
@@ 5 Conclusion
The paper presents a novel methodology that employs feature discrimination to evaluate the impact of quantization on classification accuracy, departing from the conventional reliance on quantization errors. This innovative approach provides a more direct and rational evaluation of classification performance. The study theoretically demonstrates that common binary and ternary quantization methods can improve the feature discrimination of the original data, particularly when data vectors within the same class exhibit Gaussian distributions in each dimension. These theoretical findings are supported by numerical simulations. Furthermore, the improved feature discrimination is confirmed through classification tasks conducted on both synthetic data that conform to Gaussian distributions and various real-world datasets. These datasets include convolutional features from images, spectral features from speech, and term frequency features from texts, which do not strictly follow Gaussian distributions. This underscores the robustness and generalizability of the theoretical findings.
A significant contribution of this research is its challenge to the conventional belief that larger quantization errors generally lead to worse classification performance. By doing so, it lays the groundwork for exploring more advanced quantization methods that have the potential to improve rather than degrade model performance. Additionally, it is noteworthy that binary and ternary quantization operations often result in sparse structures, a phenomenon commonly observed not only in machine learning models (Qin et al., 2020; Gholami et al., 2022) but also in biological neuron models (Haufe et al., 2010; Dasgupta et al., 2017). The findings from this research may provide valuable insights into the rationale behind the emergence of these sparse structures across various models.
@@ Appendix B Solution algorithms for Equations (8) and (9)
In this section, we present two algorithms for deriving the quantization thresholds $\beta$ that can improve feature discrimination, provided they exist as specified in Equations (8) and (9) (in Theorems 4 and 5) for binary and ternary quantization, respectively.
@@ C.2 Classification on synthetic data: KNN and SVM
C.3 Classification on synthetic data: classification accuracy vs. feature discrimination vs. quantization error
(a) Binary quantization
(b) Ternary quantization
Figure 13: KNN (Euclidean distance) classification accuracy for the binary, ternary, and original synthetic data generated with the decay parameter $\lambda=1$ and data dimension $n=1$. For comparison, the feature discrimination values and quantization errors across different thresholds $\tau$ are provided for both binary and ternary data. Comment: It is evident that higher classification accuracy tends to align with larger feature discrimination values rather than smaller quantization errors. This implies that the fluctuation in classification accuracy across $\tau$ can be more accurately captured by feature discrimination rather than by quantization errors.
C.4 Classification on real data: KNN and SVM
@@
C.5 Classification on real data: Multilayer Perceptrons (MLP) and Decision Trees
@@
C.6 Binary and Multiclass Classifications on ImageNet1000 (with 1000 classes of samples)
@@
Figure 19: Binary classification accuracy for the binary, ternary, and original data in ImageNet1000, using the classifier KNN (Euclidean distance). The parameter $\gamma$ corresponds to a quantization threshold $\tau=\gamma\cdot\eta$, where $\eta$ denotes the average magnitude of the feature elements in all feature vectors. Comment: It can be seen that for both binary and ternary quantization, there exist quantization threshold $\tau$ values that can improve the binary classification accuracy compared to the original data.

### v2

@@ 4.2.1 Setting
To validate the broad applicability of our feature discrimination analysis results, we here examine not only binary classification but also multiclass classification. Regarding the classifiers, we employ not only the fundamental KNN and SVM which mainly involve linear operations, but also more complicated classifiers involving nonlinear operations, such as the multilayer perceptron (MLP) (Rumelhart et al., 1986) and decision trees (Quinlan, 1986). For brevity, in the main text we present the classification results of YaleB, Newsgroup, and TIMIT using KNN (with Euclidean distance) and SVM, as illustrated in Figures 5 to 7. The results for other datasets, such as CIFAR10 and ImageNet1000, and other classifiers, including KNN with cosine distance, MLP and decision trees, are provided in Appendix C, specifically in Figures 14 to 20.
@@ 4.2.2 Results
Binary classification.
In Figures 5-7,14, 15 and 19, we present the binary classification results of KNN and SVM on five distinct datasets. It can be seen that there indeed exist specific ranges of quantization threshold $\tau$ values that allow both binary and ternary quantization to achieve superior, or comparable classification performance compared to the original data. Note that as illustrated in Figure 16, each data category approximately follows, rather than perfectly conforms to the Gaussian distribution assumption made in our theoretical analysis. This underscores the robustness of our theoretical results. Similar to the classification of synthetic data, the classification of real data exhibits the following performance trends. 1) When using Euclidean distance, KNN consistently obtains quantization threshold $\tau$ values that raise the classification accuracy of original data across all datasets. 2) In contrast to cosine distance, Euclidean distance enables KNN to cover a wider range of $\tau$ values that facilitate classification improvement. 3) With SVM, quantization occasionally achieves comparable performance to the original data, rather than surpassing it, as exemplified in Figure 14. 4) Ternary quantization often outperforms binary quantization by offering a broader range of threshold $\tau$ values conductive to classification improvement. The reasoning behind these results has been elaborated in our previous classification analysis of synthetic data. Overall, the consistency in classification performance between real data and synthetic data substantiates the rationality and generalizability of our theoretical findings.
KNN and SVM are well-suited for our linear discrimination analysis since they use linear operations to assess feature similarity. Other, more complex nonlinear classifiers, like MLP and decision trees, can also obtain the threshold $\tau$ values that enhance classification performance, as demonstrated in Figures 17 and 18. This capability stems from the fact that these nonlinear classifiers are grounded in linear operations, which serve to measure the linear similarity between features or model parameters, thereby aligning with our analytical framework.
@@
By enhancing feature discrimination between two classes, the performance of multiclass classification is expected to improve as well. This improvement has been observed in recent experiments (Lu et al., 2023), particularly in the challenging 1,000-class classification on deep features of ImageNet1000, as illustrated in Figure 20. In multiclass classification, we simply apply a uniform threshold to each dimension of the feature vectors. This approach conforms to our analytical framework rooted in binary classification, as classification in each dimension can be approximately considered a binary classification problem, where the values of feature elements across different classes indicate the intensity of the feature attribute at that dimension, i.e., whether it is strong or weak. The successful adaption to multiclass classification highlights the broad applicability of our theoretical findings.
@@ 5 Conclusion
The paper proposes a novel methodology that leverages feature discrimination to evaluate the influence of quantization on classification accuracy. Unlike the conventional approach, which focuses primarily on quantization errors, this feature discrimination-based analysis provides a more direct and logically sound assessment of classification performance.
The theoretical analysis reveals that common binary and ternary quantization techniques can enhance the feature discrimination of the original data, particularly when data vectors within the same class exhibit Gaussian distributions along each dimension. These theoretical insights are supported by numerical simulations. Moreover, the enhanced feature discrimination is validated through classification tasks conducted on both synthetic data adhering to Gaussian distributions and a variety of real-world datasets that do not strictly conform to Gaussian distributions, such as convolutional features extracted from images, spectral features derived from speech, and term frequency features obtained from texts. This underscores the robustness and generalizability of the theoretical findings.
A significant contribution of this research lies in its challenge to the widely accepted notion that larger quantization errors generally result in poorer classification performance. By challenging this conventional view, the study opens avenues for the exploration of more advanced quantization techniques that could potentially enhance, rather than impair, model performance. Consequently, this research is expected to drive advancements in the fields of model quantization and compression, particularly within the realms of deep neural networks (Qin et al., 2020; Gholami et al., 2022) and biological neural networks (Haufe et al., 2010; Dasgupta et al., 2017).
@@ Appendix B Solution algorithms for Equations (8) and (9)
In this section, we present two algorithms for deriving the quantization thresholds $\beta$ that can enhance feature discrimination, provided they exist as specified in Equations (8) and (9) (in Theorems 4 and 5) for binary and ternary quantization, respectively.
@@ B.2 Solution algorithm for Equation (9)
B.3 Experimental verification
Our solution algorithms are expected to produce quantization thresholds that are conductive to enhancing feature discrimination. To verify this, we evaluate the KNN classification accuracy for both binary and ternary quantization data derived from synthetic and real datasets, as described in Sections 4.1 and 4.2. In our algorithms, we set the parameter $c=10^{-3}$ for the Armijo rule and halt the iteration when $|g^{\prime}(\tau^{(k)})|_{2}<10^{-12}$ or the maximum iteration number is reached. For comparison, we also assess the performance of the conventional quantization method, which determines quantization thresholds by minimizing $\ell_{2}$-norm quantization errors (MQE) (Li et al., 2016). The experimental results are presented in Tables 1 and 2. As shown, our quantization data achieve superior classification accuracy than the conventional MQE data and the original data.
Table 1: Classification accuracy (%) of the original data, as well as the binary quantization data derived with the conventional MQE method and our method. The best results are highlighted in bold.
Dataset Original Data Binary data (MQE) Binary data (Ours)
YaleB 98.93 99.52 99.74
CIFAR10 94.08 92.58 93.54
TIMIT 92.58 94.35 94.46
Newsgroup 60.78 62.41 62.50
ImageNet 91.88 91.79 93.22
Synthetic Data 89.78 81.89 90.98
Table 2: Classification accuracy (%) of the original data, as well as the ternary quantization data derived with the conventional MQE method and our method. The best results are highlighted in bold.
Dataset Original data Ternary data (MQE) Ternary data (Ours)
YaleB 98.93 99.64 99.70
CIFAR10 94.08 96.56 96.91
TIMIT 92.58 94.10 94.15
Newsgroup 60.78 65.47 67.10
ImageNet 91.88 94.66 95.00
Synthetic Data 89.78 84.05 89.81
@@ C.2 Classification on synthetic data: KNN and SVM
C.3 Classification on real data: KNN and SVM
@@
C.4 Classification on real data: Multilayer Perceptrons (MLP) and Decision Trees
@@
C.5 Binary and Multiclass Classifications on ImageNet1000 (with 1000 classes of samples)
@@
Figure 19: Binary classification accuracy for the binary, ternary, and original data in ImageNet1000, using the classifier KNN (Euclidean distance). The parameter $\gamma$ corresponds to a quantization threshold $\tau=\gamma\cdot\eta$, where $\eta$ denotes the average magnitude of the feature elements in all feature vectors. Comment: It can be seen that for both binary and ternary quantization, there exist quantization threshold $\tau$ values that can enhance the binary classification accuracy compared to the original data.